if john's grandmother age is a perfect square, and her age equals the square of his father's age less the square of his mother's age, 40, what is the age of his grandmother?

Answer by Edwin McCravy(20056)

g = f² - 40²

factor the right side:

g = (f - 40)(f + 40)

f has to be older that 40; otherwise, (f - 40)(f + 40) would not be positive.

If f were 42 or older, the grandmother would be 164 or older, which nobody
can be.  So the father can only be 41

g = (41 - 40)(41 + 40)

g = (1)(81)

g = 81

And that is a perfect square.  So the grandmother is 81.

Edwin